Find the slope of the line passing through the points (–9, –4) and (0, –4).



A) Undefined B) 0 C) 1 D) 9


Answer by nik_consult(24)

Find the slope of the line passing through the points (–9, –4) and (0, –4).
slope, m= y2-y1 / x2-x1 ; y2=-4, y1=-4, x2=0 and x1=-9
= -4 - (-4) / 0 - (-9)
= -4 + 4 / 0 +9
= 0/ 9
= 0
ans B) 0